OpenStudy (anonymous):

subtract rad45-rad80

OpenStudy (anonymous):

\[\sqrt{45}-\sqrt{80}=\sqrt{3^2*5}-\sqrt{2^2*2^2*5}\]

OpenStudy (anonymous):

\[3\sqrt{5}-4\sqrt{5}=-\sqrt{5}\]\[\sqrt{-1}*\sqrt{5}=i \sqrt{5}\]